OpenStudy (anonymous):

how do you find the height and slant height of a cone when you are only given the radius ?

OpenStudy (anonymous):

Are you given the volume of the cone, or is what you wrote everything you were given?

OpenStudy (anonymous):

Because in order to find the slant height, you need the height. But in order to do that, we need to first find the height of the cone. The problem with that is that you don't have the volume or surface area of a cone, to determine the height. \[Slant.Height=\sqrt{r^2+h^2}\]

OpenStudy (anonymous):

I was given the radius of 5ft and lateral area of 65π ft(squared) , and i have to find the slant height, surface area and volume.

OpenStudy (anonymous):

lateral area?

OpenStudy (anonymous):

the lateral area is given which is 65π

OpenStudy (anonymous):

but how do i find the slant height and height with only having the radius and lateral area ?

OpenStudy (anonymous):

And you know \[s=\sqrt{r^2+h^2}\]

OpenStudy (anonymous):

So you can find the height now.

OpenStudy (anonymous):

Give it a try for me by substituting s with what I wrote down above and see what you get for the height.

OpenStudy (anonymous):

what does the "s" stand for?

OpenStudy (anonymous):

slant height.

OpenStudy (anonymous):

I already gave you the formula for the slant height, so you subsitute the slant height witht he formula.

OpenStudy (anonymous):

substitute* the*

OpenStudy (anonymous):

i can only substitute the radius because that's all i have. \[s = \sqrt{5^{^{2 }}} + h ^{2}\]

OpenStudy (anonymous):

i don't know what to do next.

OpenStudy (anonymous):

How would you find the lateral area of a cone?

OpenStudy (anonymous):

@wisejay_ Do you know the lateral formula of a cone? If you do, isn't it this: \[Lateral.Area=\pi\times r \times s\]

OpenStudy (anonymous):

@Azteck but that's the thing, i don't need to find the lateral are nor the radius because they are already given. I'm supposed to find the volume, height, surface area, and slant height of the cone. But i don't know how to do that i don't get it.

OpenStudy (anonymous):

That's where you're stumped. You need to work backwars. When you're given the lateral area and the radius, you have to ask yourself that if you got the lateral area, you would obviously need to know the formula for the lateral area. And when they ask to find the height and slant height, most of the time, you would start by finding the height. And when you know that the slant height formula is sqrt(r^2+h^2) Then you can substitute the "s" in the lateral area formula with sqrt(r^2+h^2).

OpenStudy (anonymous):

backwards*

OpenStudy (anonymous):

Okay, here's the formula for the lateral area of a cone. I will let L stand for the lateral area of the cone. \[L=\pi \times r \times s\] Since: \[s=\sqrt{r^2+h^2}\] YOu can substitute the "s" in the lateral area formula with the above expression: \[L=\pi\times r\times \sqrt{r^2+h^2}\]

OpenStudy (anonymous):

Now you only have one unknown variable which is h. And h is the height of the cone which is one of the things you need to find.

OpenStudy (anonymous):

@wisejay_ You there?

OpenStudy (anonymous):

yeah i'm here

OpenStudy (anonymous):

@Azteck can you help me substitute

OpenStudy (anonymous):

I just did it. Didn't you read?

OpenStudy (anonymous):

You know what L is which is 65 and you know what r is which is 5. Now rearrangle the equation to make h^2 the subject.

OpenStudy (anonymous):

rearrange*

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

So after you've done that, could you tell me what the height is.

OpenStudy (anonymous):

i got 2404.5

OpenStudy (anonymous):

That's way off the mark. Could you show me your working please?

OpenStudy (anonymous):

Ugh, i give up. I substituted everything in and i still don't get it

OpenStudy (anonymous):

Don't give up. Never give up. Maths is s tep by step process. You don't just go all out in the beginning and have nothing left at the end. You have to take it one step at a time. Not everything at once.

OpenStudy (anonymous):

Okay, first: You have this equation: \[L=\pi \times r \times \sqrt{r^2+h^2}\] Substitute your given values and you get: \[65=5\pi\sqrt{25+h^2}\]

OpenStudy (anonymous):

Now you need to make h^2 the subject in order to find h. But h^2 is inside the square root sign, so what should you do to get rid of the square root sign?

OpenStudy (anonymous):

do i have to multiply 5pie first ?

OpenStudy (anonymous):

Nope, what do you normally do to get rid of square root signs or radicals?

OpenStudy (anonymous):

i simplify, by getting two number that multiply to them ?

OpenStudy (anonymous):

First, be as simple as you can. This question is just using your basics in primary school. What would you do in this example to get rid of radicals. \[\sqrt{x}=2\]

OpenStudy (anonymous):

sorry for getting distracted

OpenStudy (anonymous):

well um find the square root ?

OpenStudy (anonymous):

Read what I wrote again. I asked you what you should do to get RID of the square root...

OpenStudy (anonymous):

It seems that you don't normally focus on what you need to do, except procrastinating. That's what I used to do and I was academically ranked near last in Year 7-8. Then I realised I was a terrible student who would often procrastinate. Don't follow that route to procrastination. It would lead you to nowhere but downhill...

OpenStudy (anonymous):

What grade are you in ? @Azteck

OpenStudy (anonymous):

I think we're diverting to another topic, can we continue on this question and actually get though it please?

OpenStudy (anonymous):

Okay, sorry.

OpenStudy (anonymous):

Um square root both sides ?

OpenStudy (anonymous):

so you're saying we need to square root, a square root?

OpenStudy (anonymous):

Wouldn't that complicate things even more? \[\sqrt{\sqrt{x}}=\sqrt{2}\]

OpenStudy (anonymous):

LMAO no, i mean square both sides ?

OpenStudy (anonymous):

Yes, that's what we should do with your question. \[65=5\pi\sqrt{r^2+h^2}\] \[65^2=25\pi^2(r^2+h^2)\]

OpenStudy (anonymous):

Now expand/distribute the RHS (Right Hand Side for me please?

OpenStudy (anonymous):

(Right Hand Side)*

OpenStudy (anonymous):

\[4225=246.14(25+h ^{2}) \]

OpenStudy (anonymous):

is that wrong ?

OpenStudy (anonymous):

i did it on a calculator

OpenStudy (anonymous):

Do you know what Expanding/Distributing is?

OpenStudy (anonymous):

Never mind that, you're meant to move the term outside of those brackets to the LHS. \[\frac{65^2}{25\pi r^2}=25+h^2\]

OpenStudy (anonymous):

Now make h^2 the subject. Do you know what that means when I say make h^2 the SUBJECT?

OpenStudy (anonymous):

sadly i don't @Azteck

OpenStudy (anonymous):

You need to place h^2 on it's own on one side. If you make x the subject in this example: \[x+y=2\] You would move the y to the other side: \[x=2-y\] That's making x the subject.

OpenStudy (anonymous):

Now make h^2 the subject for me please.